Please help with all 3 parts!

Salvadores Manufacturing builds and sells snowboards, skis and poles. The sales price and variable cost for each follows:

Product Selling Price
per Unit
Variable Cost
per Unit
Snowboards $340 $160
Skis $380 $210
Poles   $50   $30

a. Their sales mix is reflected in the ratio 7:3:2. What is the overall unit contribution margin for Salvadores with their current product mix?

Product Selling Price
per Unit
Variable Cost
per Unit
Snowboards $340           $190          
Skis $390           $220          
Poles $50           $30          

b. Their sales mix is reflected in the ratio 8:3:2. If annual fixed costs shared by the three products are $245,000, how many units of each product will need to be sold in order for Salvadores to break even?

Product Ratio (mix) Break-even per
composite unit
Number of Units
per product
Snowboards 8
Skis 3
Poles 2

Overall Unit Contribution Margin $

Salvador Manufacturing builds and sells snowboards, skis and poles. The sales price and variable cost for each follows:

Product Selling Price
per Unit
Variable Cost
per Unit
Snowboards $340          $150         
Skis $420          $210         
Poles $60          $10         

c. Their sales mix is reflected in the ratio 7:3:1. If annual fixed costs shared by the three products are $241,200. Determine the break-even point in sales dollars.

Break-even point $
Answer #1

(a.) Calculation of the overall unit contributon margin for salvadores with current product mix-

Snow boards Skis Poles Total 340 380 50 160 210 30 a b с d e Selling price per unit variable cost per unit Contribution margi

(b.) Calculation of BEP -

BEP (in units) = Total fixed cost/(weighted average contribution margin per unit)

Skis Poles Total Snow boards 340 190 390 501 30 220 a b с d е Selling price per unit variable cost per unit Contribution marg

BEP (in unit) composite = 245000/134.62

= 1820 units

Snowboard units sold for break even = composite unit*sales mix of snowboard

= 1820*8/13

= 1120 units

Skies unit for break even = 1820*3/13

= 420 units

poles unit for break even = 1820*2/13

= 280 units

(c) Calculation of BEP in sales $ -

Snow boards Skis Poles Total a 340 420 60 b 150 210 10 С Selling price per unit variable cost per unit Contribution margin pe

BEP (in units) composite = 241200/182.73

= 1320 units

units of snoboard = 1320*7/11 = 840 units

units of skies = 1320*3/11 = 360 units

units of poles = 1320*1/11 = 120 units

Total sales in $ = units of snowboard * sales price per unit of snowbaord + units of skies * sales price per unit of skies + units of poles * sales price per unit of poles

= 840*340 + 360*420 + 120*60

= 285600 + 151200 + 7200

= $ 444000

Please check with your answer and let me know.